# MA-GTS: A Multi-Agent Framework for Solving Complex Graph Problems in Real-World Applications

### v1

@@ 4 G-REAL
Existing datasets for evaluating LLMs’ understanding and reasoning on graph-structured data are explicitly constructed. However, real-world graph-theoretic problems often involve rich textual semantic information and implicitly structured representations. To assess the performance of the MA-GTS framework on practical problems, we introduce G-REAL, a dataset that captures real-world graph problems. This dataset comprises three commonly encountered graph-theoretic challenges: (1) the optimization of logistics and delivery routes, (2) wireless network channel allocation, and (3) network monitoring optimization. These correspond to three fundamental graph problems: the Traveling Salesman Problem (TSP), the Minimum Graph Coloring Problem, and the Minimum Vertex Cover Problem, respectivelyHoffman et al. (2013); Jensen and Toft (2011); Hochbaum (1982). The composition of G-REAL can be seen briefly in Figure 2. In this section, we provide a detailed description of the dataset’s composition and construction methodology. More detail about G-REAL in Appendix C.
G-REAL GraCoRe NLGraph TSP Coloring Vetex Cover TSP Shortest Path Cycle #Graph 900 900 900 360 380 1150 Node Range 8 - 25 8 - 25 8 - 25 8 - 25 5 - 20 5 - 15 Real-World Problem Text Noise
@@
G-REAL GraCoRe NLGraph
Model Method TSP Coloring Vetex Cover TSP Shortest Path Cycle
Direct 11.8% 80.1% 68.7% 79.7% 100.0% 98.9%
o3-mini CoT 12.9% 83.1% 72.8% 80.0% 98.4% 98.9%
Direct 2.5% 23.4% 0.3% 1.1% 27.3% 50.9%
GPT-4o-mini CoT 3.1% 25.1% 0.0% 1.1% 27.6% 51.1%
Direct 0.1% 0.7% 2.5% 1.9% 30.5% 50.0%
GPT-3.5 CoT 2.1% 7.6% 4.8% 1.6% 34.7% 49.9%
Direct 0.6% 16.2% 17.4% 3.8% 22.1% 49.6%
Qwen2.5-7b CoT 0.6% 8.8% 8.5% 3.0% 27.3% 52.7%
Direct 3.6% 10.1% 7.2% 0.3% 12.6% 53.7%
Llama 3 -7b CoT 4.1% 14.3% 6.7% 0.3% 19.4% 50.9%
MA-GTS(GPT-4o-mini) Multi-Agent 94.9%($\uparrow$82%) 94.5%($\uparrow$11.4%) 93.2%($\uparrow$20.4%) 96.9%($\uparrow$14.9%) 97.8%($\downarrow$2.2%) 98.9%(0%)
@@ 4.3 Evaluation
[…]
@@ 5.1 Datasets
To evaluate the reasoning capabilities of the MA-GTS framework across various graph-theoretic problem types, complexities, and domains, we used the G-REAL dataset alongside two benchmark datasets, GraCoReYuan et al. (2025) and NLGraphWang et al. (2024), covering six distinct graph-theoretic tasks. We selected three sub-tasks for evaluation: the TSP, shortest path problem, and Cycle problem in GraCoRe and NLGraph. Notably, both GraCoRe and G-REAL include TSP instances; however, the G-REAL TSP is more complex and reflects real-world scenarios with implicit graph structure data. By comparing performance on these two TSP instances, we assess the model’s ability to handle more intricate problems. The simpler tasks in NLGraph evaluate the generalization and robustness of the MA-GTS framework. A summary of the differences between these datasets is provided in Table 1.
@@ 5.2 Baselines and Foundation Model
We compared three of OpenAI’s latest closed-source models: o3-mini, GPT-4o-mini, and GPT-3.5Achiam et al. (2023). Additionally, we evaluated two of the most recent open-source models: Llama3-7bTouvron et al. (2023) and Qwen2.5-7bBai et al. (2023). For the evaluation methodology, we adopted both direct inference and CoT reasoning approaches. For the foundation model, we selected the GPT-4o-mini model. Regarding the final test results, for each task, we used the accuracy of the final computed solution as the primary evaluation metric. More details about models in Appendix B.
@@ 6 Results and Analysis
In this section, we evaluate the performance of our framework against other LLMs on graph theory problems, with results presented in Table 2. MA-GTS outperforms all baselines, achieving state-of-the-art results and matching the performance of the leading o3-mini model on simpler problems. We also assess the MA-GTS framework from multiple perspectives.
TSP Coloring VertexCover
Inp.Tokens(k) Out.Tokens(k) Price(\$) Inp.Tokens(k) Out.Tokens(k) Price(\$) Inp.Tokens(k) Out.Tokens(k) Price(\$)
o3-mini 2.31 4.98 0.0244 0.69 6.83 0.0309 0.6 9.73 0.0443
MA-GTS(GPT-4o-mini) 13.32 4.56($\downarrow$8.4%) 0.0047($\downarrow$80.7%) 6.79 2.57($\downarrow$62.4%) 0.0025($\downarrow$91.9%) 6.39 2.31($\downarrow$76.2%) 0.0023($\downarrow$94.8%)
@@ 6.1 Performance on real-world problems
As shown in Table 2, G-REAL provides three real-world graph theory problems, with the TSP being the most complex. Based on the results from these three problems, MA-GTS demonstrates superior performance, achieving an accuracy rate exceeding 90% across all tests. Notably, in the case of the TSP, MA-GTS outperforms the o3-mini model by 82%. Furthermore, when compared to the GPT-4o-mini model, MA-GTS significantly improves its performance from 3.1% to 94.9%, marking a substantial increase. This clearly underscores the effectiveness of our framework. Additionally, it is evident that, aside from the o3-mini model, other models exhibit subpar performance on the G-REAL dataset. It is particularly interesting that the performance gap between the two open-source and two closed-source models is minimal, suggesting that the complexity of the problems may lead to a consistent decline in performance, an issue that warrants further investigation. Overall, MA-GTS stands out for its advanced capabilities and generalization when handling complex graph theory problems.
@@ 6.2 Performance on simple problem
Table 2 shows that for simpler graph theory problems, such as the Shortest Path and Cycle problems from the NLGraph dataset, the o3-mini model performs exceptionally well, with MA-GTS also showing strong results. Specifically, for the Shortest Path problem, the gap between MA-GTS and o3-mini is just 2.2%, and MA-GTS performs equally well on the Cycle problem. In contrast, other models perform less satisfactorily. The MA-GTS framework, based on the GPT-4o-mini model, significantly enhances the accuracy of the 4o model, bringing it on par with the o3-mini. Overall, MA-GTS demonstrates excellent performance across diverse textual descriptions and graph structures, highlighting its remarkable generalization capabilities.
@@ 6.3 G-REAL effectiveness analysis
To evaluate the performance of LLMs and MA-GTS on real-world graph theory problems, we constructed the G-REAL dataset. As shown in Table 2, the performance of existing LLMs on the G-REAL dataset is suboptimal. To validate the effectiveness of this dataset, we compared it with the TSP problem from the GraCoRe Benchmark, testing problems with node sizes ranging from 8 to 25, consistent with the scale of G-REAL. From this comparison, we observe that on the G-REAL dataset, which includes text complexity, added text noise, and node name shuffling, the o3-mini model performs poorly, with its accuracy dropping from 79.7% in GraCoRe to 11.8%. In contrast, the MA-GTS framework appears unaffected by the complexities of real-world graph theory problems, maintaining performance above 90%. This result indirectly supports the validity of the G-REAL dataset and demonstrates the stability of the MA-GTS framework.
Figure 3: Performance of different problems across varying node numbers, showing results for MA-GTS and o3-mini.
G-REAL TSP Coloring Vetex Cover Average error rate GPT-4o-mini(Tool use) 30.8% 39.0% 4.6% 75.0% w/o IEL 12.5% 42.2% 14.6% 19.4% w/o KIL 7.8% 37.1% 12.8% 1.0% w/o AEL 4.6% 32.1% 7.4% 3.2% MA-GTS(GPT-4o-mini) 94.9% 94.5% 93.2% 0.5%
Table 4: Testing Results and Error Analysis of Ablation Experiments for Each Layer of MA-GTS. "Tool use" refers to the utilization of only the algorithm library we have constructed.
@@ 6.4 Impact of Node Size
To evaluate the impact of node scale on LLMs in complex graph theory problems, we tested the performance of MA-GTS and the o3-mini model on four complex graph problem datasets, with node sizes ranging from 8 to 25. The results, shown in Figure 3, clearly demonstrate that as the number of nodes increases, the performance of the o3-mini model deteriorates, particularly in the TSP problem from G-REAL. For node sizes greater than 20, the o3-mini model is unable to produce correct answers. In contrast, under the MA-GTS framework, the effect of node size is less pronounced. Even with more than 20 nodes, MA-GTS maintains high prediction accuracy and stability. This highlights both the effectiveness and superiority of our framework.
@@ 6.5 Cost Analysis
Since MA-GTS requires multiple agent calls to model APIs for inference, cost considerations arise. To address this, we compared the inference costs of MA-GTS based on the GPT-4o-mini model with the o3-mini model, as shown in Table 3. Surprisingly, MA-GTS incurs significantly lower costs than the o3-mini model. The o3-mini model, in contrast, has hidden reasoning tokens during inference, leading to long, concealed reasoning processes even in direct inference scenarios. As shown in the table, the inference cost of MA-GTS is about one-tenth to one-twentieth of the o3-mini model, requiring far fewer inference tokens. Moreover, MA-GTS achieves far better results than o3-mini, demonstrating its high cost-effectiveness in delivering more accurate outcomes at a lower cost.
@@ 6.6 Ablations Studies and Analyses
To validate the effectiveness of each layer in MA-GTS, we conducted ablation experiments, with results shown in Table 4. The table demonstrates that each layer is crucial, and removing any layer significantly affects the final results. Although the IEL layer has the smallest impact on accuracy, its absence leads to a substantial increase in error rate (19%), highlighting its role in maintaining stability. The absence of the AEL layer results in the greatest accuracy loss. Even when a module is removed, MA-GTS still improves the accuracy of the base model, validating the framework’s effectiveness. Additionally, when inference is performed using only the GPT-4o-mini model with the constructed algorithm library, accuracy improves, but the error rate remains high (75%). For graph sizes larger than 10 nodes, the model struggles to correctly invoke algorithms, further demonstrating the robustness and generalizability of MA-GTS.
@@ 7 Conclusion
This paper introduces MA-GTS, a Multi-Agent Framework for solving real-world graph theory problems, validated using the G-REAL dataset. Performance comparisons across various LLM models show that MA-GTS achieves high accuracy, stability, and cost-effectiveness, excelling in both complex and simpler graph problems. With accuracy consistently above 90%, MA-GTS outperforms traditional LLM approaches, maintaining stability across different problem scales and being well-suited for larger graphs. Future work will focus on scaling to even larger problems and improving cost-efficiency.
@@ A.3 Prompt Templates
In this section, I will introduce the prompts for each agent, which will be displayed in Figures 6 to 11.
@@ Appendix B Details on baseline models
We evaluated 5 of the latest LLMs, including OpenAI o3-mini reasoning model, launched on January 31, 2025. Table 5 presents more details on the models and their versions.
@@ Appendix C Details on G-REAL
Existing graph theory benchmarks do not align with real-world scenarios. To better evaluate the ability of MA-GTS in solving graph theory problems in practical contexts and to test the performance gap between LLMs on structured textual graph data and implicit representations, we constructed the G-REAL dataset. This dataset contains three common real-world problems, with detailed information provided in the G-REAL section. It generates problem graphs of varying scales by randomly encoding node names and structures, with the naming conventions and sample problems illustrated in Figures 12 to 15.

### v2

@@ 4 G-REAL
Existing datasets for evaluating LLMs’ understanding and reasoning on graph-structured data are explicitly constructed. However, real-world graph-theoretic problems often involve rich textual semantic information and implicitly structured representations. To assess the performance of the MA-GTS framework on practical problems, we introduce G-REAL, a dataset that captures real-world graph problems. This dataset comprises four commonly encountered graph-theoretic challenges: the Traveling Salesman Problem (TSP), the Minimum Graph Coloring Problem, the Minimum Vertex Cover Problem and the Shortest Path Problem, respectively hoffman2013traveling; jensen2011graph; hochbaum1982approximation. They correspond to four common problems in real-world scenarios, namely the Delivery Logistics Problem, the Wireless Channel Allocation Problem, the Network Monitoring Problem, and the Target Navigation Problem. The composition of G-REAL can be seen briefly in Figure 3. In this section, we provide a detailed description of the dataset’s composition and construction methodology. More detail about G-REAL in Appendix C.
[…]
@@
[…]
@@ 4.3 Evaluation
For the TSP, Minimum Graph Coloring, Minimum Vertex Cover, and Shortest Path problems, the output includes both selected nodes and the final solution, requiring dual evaluation. To fully assess LLMs’ graph reasoning, both output types are used as evaluation metrics. The model’s performance is measured by verifying the accuracy of both the selected node set and the computed solution. The methodology for calculating the final accuracy is as follows: $\text{$ACC_{ALL}$}=0.5\cdot ACC_{\text{nodes}}+0.5\cdot ACC_{\text{result}},$ where $ACC_{\text{nodes}}$ and $ACC_{\text{result}}$ represent the accuracy of the node set and the predicted values, respectively, with a value of 1 for correct predictions and 0 for incorrect ones.
@@ 5.1 Datasets
To evaluate the reasoning capabilities of the MA-GTS framework across various graph-theoretic problem types, complexities, and domains, we used the G-REAL dataset alongside two benchmark datasets, GraCoRe yuan2025gracore and NLGraph NLgraph-6, covering seven distinct graph-theoretic tasks. We selected three sub-tasks for evaluation: the TSP, shortest path problem, and Cycle problem in GraCoRe and NLGraph. Notably, both GraCoRe and G-REAL include TSP instances, both NLGraph and G-REAL include Shortest Path instances; however, the G-REAL TSP and Shortest Path is more complex and reflects real-world scenarios with implicit graph structure data. By comparing performance on these two types instances, we assess the model’s ability to handle more intricate problems. The simpler tasks in NLGraph evaluate the generalization and robustness of MA-GTS. A summary of the differences between these datasets is provided in Table 1.
@@ 5.2 Baselines and Foundation Model
We compared three of OpenAI’s latest closed-source models: o3-mini, GPT-4o-mini, and GPT-3.5 gpt4. Additionally, we evaluated three of the most recent open-source models: Llama3-7b llama, Qwen2.5-7b qwen and Deepseek-V3-660B liu2024deepseek. For the evaluation methodology, we adopted both direct inference and CoT reasoning approaches. For the foundation model, we selected the GPT-4o-mini and Deepseek-V3-660B model, they respectively represent some of the more advanced open-source and closed-source models. Furthermore, we conducted a comparative analysis of the performance of OWL owl2025 and GraphTeam li2024graphteam, which respectively represent a general-purpose multi-agent framework and a graph-theoretic multi-agent framework. Regarding the final test results, for each task, we used the accuracy of the final computed solution as the primary evaluation metric. More details about models in Appendix B.
@@ 6 Results and Analysis
We evaluate the performance of our framework against other LLMs on graph theory problems, with results presented in Table 2. MA-GTS outperforms all baselines, achieving state-of-the-art results and matching the performance of the leading o3-mini model on simpler problems. We also assess the MA-GTS framework from multiple perspectives.
[…]
@@ 6.1 Performance on real-world problems
As shown in Table 2, G-REAL provides four real-world graph theory problems, with TSP being the most complex. Based on the results from these problems, MA-GTS demonstrates superior performance, achieving an accuracy rate exceeding 90% across all tests. Notably, in the case of the TSP, MA-GTS outperforms the o3-mini model by 82%. Even when built upon the open-source DeepSeek model, MA-GTS still achieves strong performance. Furthermore, when compared to the GPT-4o-mini model, MA-GTS significantly improves its performance from 3.1% to 94.9%, marking a substantial increase. This clearly underscores the effectiveness of our framework. Additionally, it is evident that, aside from the o3-mini model, other models exhibit subpar performance on the G-REAL dataset. It is particularly interesting that the performance gap between the two open-source and two closed-source models is minimal, suggesting that the complexity of the problems may lead to a consistent decline in performance, an issue that warrants further investigation. Overall, MA-GTS stands out for its advanced capabilities and generalization when handling complex graph theory problems.
@@ 6.2 Performance on simple problem
Table 2 shows that for simpler graph theory problems, such as the Shortest Path and Cycle problems from the NLGraph dataset, the o3-mini model performs exceptionally well, with MA-GTS also showing strong results. Specifically, for Shortest Path problem, the gap between MA-GTS and o3-mini is just 2.2%, and MA-GTS performs equally well on the Cycle problem. In contrast, other models perform less satisfactorily. The MA-GTS framework, based on the GPT-4o-mini model, significantly enhances the accuracy of the 4o model, bringing it on par with the o3-mini. Overall, MA-GTS demonstrates excellent performance across diverse textual descriptions and graph structures, highlighting its remarkable generalization capabilities.
@@ 6.3 G-REAL effectiveness analysis
To evaluate the performance of LLMs and MA-GTS on real-world graph theory problems, we constructed the G-REAL dataset. As shown in Table 2, the performance of existing LLMs on the G-REAL dataset is suboptimal. To validate the effectiveness of this dataset, we compared it with the TSP problem from the GraCoRe Benchmark, testing problems with node sizes ranging from 8 to 25, consistent with the scale of G-REAL. From this comparison, we observe that on the G-REAL dataset, which includes text complexity, added text noise, and node name shuffling, the o3-mini model performs poorly, with its accuracy dropping from 79.7% in GraCoRe to 11.8%. In contrast, the MA-GTS framework appears unaffected by the complexities of real-world graph theory problems, maintaining performance above 90%. This result indirectly supports the validity of the G-REAL dataset and shows the stability of the MA-GTS.
Figure 4: Performance of different problems across varying node numbers (MA-GTS v.s. o3-mini).
@@ 6.4 Impact of Node Size
To evaluate the impact of node scale on LLMs in complex graph theory problems, we tested the performance of MA-GTS and the o3-mini model on four complex graph problem datasets, with node sizes ranging from 8 to 25. The results, shown in Figure 4, clearly demonstrate that as the number of nodes increases, the performance of the o3-mini model deteriorates, particularly in the TSP problem from G-REAL. For node sizes greater than 20, the o3-mini model is unable to produce correct answers. In contrast, under the MA-GTS, the effect of node size is less pronounced. Even with more than 20 nodes, MA-GTS maintains high prediction accuracy and stability. It highlights both the effectiveness and superiority of MA-GTS. Performance of MA-GTS on larger node scales is discussed in the Appendix D.
@@ 6.5 Cost Analysis
Since MA-GTS requires multiple agent calls to model APIs for inference, cost considerations arise. To address this, we compared the inference costs of MA-GTS based on the GPT-4o-mini model with the o3-mini model, as shown in Table 3. Surprisingly, MA-GTS incurs significantly lower costs than the o3-mini model. The o3-mini model, in contrast, has hidden reasoning tokens during inference, leading to long, concealed reasoning processes even in direct inference scenarios. As shown in the table, the inference cost of MA-GTS is about one-tenth to one-twentieth of the o3-mini model, requiring far fewer inference tokens. Moreover, MA-GTS achieves far better results than o3-mini, demonstrating its high cost-effectiveness in delivering more accurate outcomes at a lower cost. Runtime efficiency is discussed in the Appendix D.
[…]
Table 4: Ablation Experiments for Each Layer of MA-GTS ("Tool use" refers to the utilization of only the algorithm library we have constructed).
@@ 6.6 Ablations Studies and Analyses
To validate the effectiveness of each layer in MA-GTS, we conducted ablation experiments, with results shown in Table 4. It demonstrates that each layer is crucial, and removing any layer significantly affects the final results. Although the IEL layer has the smallest impact on accuracy, its absence leads to a substantial increase in error rate (19%), highlighting its role in maintaining stability. The absence of the AEL layer results in the greatest accuracy loss. Even when a module is removed, MA-GTS still improves the accuracy of the base model, validating the framework’s effectiveness. Additionally, when inference is performed using only the GPT-4o-mini model with the constructed algorithm library, accuracy improves, but the error rate remains high (75%). For graph sizes larger than 10 nodes, the model struggles to correctly invoke algorithms, further demonstrating the robustness and generalizability of MA-GTS.
[…]
Table 5: Sensitivity analysis of multi-round inference.
6.7 Sensitivity Studies of MA-GTS
We have conducted an additional sensitivity experiment for MA-GTS. For each graph dataset size, we randomly selected 5 problems and queried MA-GTS 5 times per problem to test the stability and sensitivity of the framework. As shown in Table 5, the results demonstrate that MA-GTS is highly stable and there is no significant performance fluctuation across repeated queries.
@@ 7 Conclusion
We introduces MA-GTS, a Multi-Agent Framework for solving real-world graph theory problems, validated using the G-REAL dataset. Performance comparisons across various LLMs show that MA-GTS achieves high accuracy, stability, and cost-effectiveness, excelling in both complex and simpler graph problems. With accuracy consistently above 90%, MA-GTS outperforms existing methods, maintaining stability across different problem scales and being well-suited for larger graphs. Future work will focus on scaling to even larger problems and improving cost-efficiency.
@@
Acknowledgments
This work is supported by the National Science Foundation of China (U22B2059, 62276083), the Major Key Project of PCL (Grant No. PCL2024A08) and the 5G Application Innovation Joint Research Institute’s Project (A003).
@@ A.3 Prompt Templates
In this section, I will introduce the prompts for each agent, which will be displayed in Figures 7 to 12.
@@ Appendix B Details on baseline models
We evaluated 6 of the latest LLMs, including OpenAI o3-mini reasoning model, launched on January 31, 2025 and the latest open-source model, DeepSeek-V3. Table 9 presents more details on the models and their versions.
@@ Appendix C Details on G-REAL
Existing graph theory benchmarks do not align with real-world scenarios. To better evaluate the ability of MA-GTS in solving graph theory problems in practical contexts and to test the performance gap between LLMs on structured textual graph data and implicit representations, we constructed the G-REAL dataset. This dataset contains three common real-world problems, with detailed information provided in the G-REAL section. It generates problem graphs of varying scales by randomly encoding node names and structures, with the naming conventions and sample problems illustrated in Figures 13 to 16.
Appendix D More experimental analysis
D.1 Large-scale node analysis
Why 8-25 enough: In our work, we chose to focus on graphs with 8–25 nodes, primarily due to the complexity and reasoning difficulty posed by realistic tasks. Unlike large-scale but structurally explicit graphs, the G-REAL dataset introduces substantial textual noise, implicit graph structures, and randomly named nodes. These factors make the problem setting significantly closer to real-world semantic reasoning scenarios and increase the overall problem difficulty. This differs from existing benchmarks, which typically construct graph-structured data using explicitly defined and concise textual descriptions.
More experiments: We have extended our experiments to include larger graph sizes. As shown in Table 6, we tested the TSP and Graph Coloring problems with 25, 30, 35, and 40-node graphs, with 5 instances evaluated for each size. The results show that even on larger graphs, our framework maintains high accuracy and stability. We plan to include more experiments on even larger graph sizes in future versions of the paper to further validate the scalability of our approach.
25 30 35 40 TSP 0.8 0.6 0.8 0.8 Coloring 1 0.6 1 0.6
Table 6: This table shows the results of experiments conducted on TSP and Coloring problems with extended graph sizes of 25, 30, 35, and 40 nodes. For each size, 5 problem instances were tested.
D.2 Runtime analysis
G-REAL focuses on graph reasoning under complex semantic conditions, which more closely resemble real-world user scenarios. These tasks often contain intricate semantic information and irrelevant noise, posing significant challenges for LLMs. For instance, the Delivery Logistics Problem in G-REAL is a TSP instance. In contrast to existing TSP benchmarks—where nodes are ordered, connections are explicitly stated, and the problem type is clearly defined—G-REAL requires the model to infer all of this information from natural language. This increases the difficulty of graph construction and makes reasoning more error.
We conducted a supplementary evaluation on the G-REAL-TSP task by randomly selecting five graphs with 15-node scales (Table 7). We measured the average solution time per problem instance, including task decomposition, tool invocation, and result verification. Compared to existing multi-agent frameworks, MA-GTS demonstrates a clear advantage in time efficiency. These results highlight that our framework is capable of maintaining high accuracy while keeping inference time relatively low, further validating its practical applicability.
MA-GTS Graphteam OWL Time use (s) 148.48 251.34 139.39 ACC (%) 100 0 0
Table 7: The table presents the results of testing on the G-Real-TSP problem using 5 randomly selected graphs with 15 nodes each.Base model is GPT-4o-mini.
Appendix E G-REAL Details
G-REAL introduces two main innovations over existing graph reasoning benchmarks. First, it situates graph problems in realistic task contexts with background noise, where node names are randomized and both node and edge descriptions contain distracting language. This design substantially increases the difficulty of structure extraction and semantic understanding, in contrast to benchmarks such as GraphArena that rely on clean, explicit inputs. Second, G-REAL effectively mitigates data contamination by ensuring that randomized identifiers and artificial noise prevent test samples from matching pretraining corpora. As summarized in Table 8, these features make G-REAL more representative of real-world tasks, more robust against leakage, and more focused on challenging semantic reasoning—aligning it closely with the practical difficulties faced by LLMs.
[…]
Table 8: Comparison with mainstream benchmarks.
@@ Appendix E G-REAL Details
Deepseek-V3 DeepSeek-V3-0324-660B https://huggingface.co/deepseek-ai/DeepSeek-V3-0324